Determine the total capacitance if a 30  pF  capacitor is connected in series with a 0.03 μF capacitor.     Capacitance =


1
Expert's answer
2022-02-28T13:52:49-0500

Gives

"C_1=30pF 
\\C_2=30pF"

Now "C_1" And C"_2" Is series combination

"C_{net}=\frac{C_1C_2}{C_1+C_2}"

"C{net}=\frac{30\times30}{30+30}pF=15pF"